OpenStudy (anonymous):

2x 3+ 4x 2 + 8x a. 2x3 + 4x2 + 8x b. x(2x2 + 4x + 8) c. 2x(x2 + 2x + 4) d. 2x(x + 2)(x + 4) 2 points QUESTION 2 1. Complete. z2 + 9z – 90 = (z – 6)(z + ) a. –15 b. 90 c. –9 d. 15 2 points QUESTION 3 1. Factor the expression. w2 + 18w + 77 a. (w – 7)(w – 11) b. (w + 7)(w + 11) c. (w + 1)(w + 77) d. (w – 7)(w + 11) 2 points QUESTION 4 1. Factor the expression. d2 + 13d + 36 a. (d – 9)(d – 4) b. (d + 9)(d + 4) c. (d – 9)(d + 4) d. (d + 9)(d – 4)

OpenStudy (janu16):

What is the largest number that divides evenly into 2x^3,4x^2,8x^2?

OpenStudy (janu16):

What is the highest degree of xx that divides evenly into 2x^3,4x^2,8x^2?

OpenStudy (janu16):

answer those 2 questions first for number 1

OpenStudy (janu16):

@brandel26

OpenStudy (anonymous):

into

OpenStudy (anonymous):

im really bad at algebra

OpenStudy (janu16):

what?

OpenStudy (janu16):

so we know that the GCF is 2x so now 1)Put the GCF as the first term 2)Then, in parentheses, divide each term by the GCF

OpenStudy (janu16):

\[2x(\frac{ 2x^3 }{ 2x }+\frac{ 4x^2 }{ 2x }+\frac{ 8x }{ 2x })\]

OpenStudy (janu16):

divide everything in parenthesis and what do you get?

OpenStudy (anonymous):

7x 6?

OpenStudy (janu16):

no what is 2 divide by 2?

OpenStudy (janu16):

2x^3 divided by 2x?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1x 4

OpenStudy (janu16):

no great so x^2 now what is 4x^2 divided by 2x?

OpenStudy (janu16):

and lastly what is 8x divided by 2x?

OpenStudy (anonymous):

4x 2

OpenStudy (janu16):

so 4 divided by 2 is 2 and 8 divided by 2 is 4 corrrect?

OpenStudy (anonymous):

yes

OpenStudy (janu16):

you are just divind the top and bottom number which is easy but so we have 2x(x^2+2x+4) and we are done for number 1

OpenStudy (janu16):

so what choice is number 1?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

c

OpenStudy (janu16):

no its c

OpenStudy (janu16):

yup

OpenStudy (janu16):

for number 2 whats after z+ cause you didnt write it

OpenStudy (anonymous):

its blsnk

OpenStudy (janu16):

so its (z+)?

OpenStudy (anonymous):

on the test theres a blank box

OpenStudy (janu16):

are you sure? do you have a picture?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

yes im sure

OpenStudy (janu16):

i got the answer but doesnt match your choices so i dont know what to do is there only z +

OpenStudy (anonymous):

im gonna show you

OpenStudy (anonymous):

i cant screenshot it

OpenStudy (janu16):

is tht flvs?

OpenStudy (janu16):

@brandel26

OpenStudy (anonymous):

got it. Also thanks for the help

OpenStudy (janu16):

do you need other 2?

OpenStudy (anonymous):

yes please

OpenStudy (janu16):

w2 + 18w + 77 so for number 3 which 2 numbers add and multiplies to 18 and 77?

OpenStudy (janu16):

what two SAME numbers equals 18 when you add and equals 77 when you multiply?

OpenStudy (janu16):

@brandel26

OpenStudy (anonymous):

7 and 11?

OpenStudy (janu16):

yess so now we are gonna put tht in pranthesis

OpenStudy (janu16):

(w+7)(w+11) thats the answer good job so what choice?

OpenStudy (anonymous):

b

OpenStudy (janu16):

correct

OpenStudy (anonymous):

alright

OpenStudy (janu16):

so for the last one we are gonna do the same thing. d2 + 13d + 36 what two SAME numbers equals 13 when you add and equals 36 when you multiply?

OpenStudy (anonymous):

9 and 4

OpenStudy (janu16):

so now we are gonna put thtat in pranthsis (d+9)(d+4)

OpenStudy (janu16):

what chocie?

OpenStudy (anonymous):

b

OpenStudy (janu16):

correct and that was the last one right?

OpenStudy (janu16):

i mean last problem

OpenStudy (anonymous):

i have a more but i cant make you do all lol

OpenStudy (janu16):

just put that in new thread and put 1 for each so it doesnt lag

OpenStudy (anonymous):

reply too this?

OpenStudy (janu16):

no go to question box

OpenStudy (janu16):

and post one question

OpenStudy (anonymous):

ok

OpenStudy (janu16):

its on the top left corner on computer